# RepoExec: Evaluate Code Generation with a Repository-Level Executable Benchmark

### v2

@@ 5 Experiment
5.1 Evaluation metrics
We use the well-known Pass@k Chen et al. (2021) to measure the functional correctness of generation outputs. Futhermore, to assess the models’ ability to utilize provided dependencies in accordance with human intent, we introduce the Dependency Invocation Rate (DIR). This metric represents the percentage of invoked dependencies out of the total number of dependencies provided. Let $D_{g}$ denote the set of identifiers in the generated output, and $D_{s}$ denote the set of provided dependencies extracted from the solution. The Dependency Invocation Rate (DIR) is calculated as follows:
$\text{DIR}=\dfrac{|D_{g}\cap D_{s}|}{|D_{s}|}$
@@ 5.3 Generation with Multi-round Debugging
In this section, we examine the models’ self-refinement capabilities in enhancing generation performance. We provide the models with error output logs and ask them to fix the errors. We experiment with WizardCoder, GPT3.5, and CodeLlama-13b-Python. The number of rounds to debug is set to 3 and the input template is presented in Appendix F. In this experiment, we employ a greedy search algorithm to generate only a single output for each problem.
@@ 5.4 Instruction-tuning with Code Dependencies
While the multi-round debugging experiment has demonstrated effectiveness in leveraging given dependencies to provide correct solutions (Section 5.3), it requires a strong model to generate good test cases and can be time-consuming due to the repeated generation and execution of code and test cases. To address these challenges, we propose an instruction-tuning dataset for fine-tuning base LLMs. We collected the 1,555 most-starred repositories from 2018 onward and extracted functions with their corresponding dependencies, following the procedure outlined in Section 3.2. We obtained 154,818 functions, of which 57,746 samples include docstrings. From this set, we utilized 50K samples with docstrings and applied instruction prompts, while 80K samples adhered to the raw code format (Full context). Recognizing the potential of the Small-context format, we allocated the remaining 20K samples to follow this structure. We fine-tuned StarCoder, StarCoder2, and CodeLlama-13b-Python models for 5 epochs with LoRa Hu et al. (2021) and used 10% of the training data as the validation set to select the best checkpoint.
Table 5 illustrates the efficacy of our training data. All 3 models demonstrate improvements in both Pass@1 and DIR after instruction tuning. Specifically, there is a slight increase of around 1% in Pass@k for all models, while DIR shows a significant improvement, reaching the highest scores (over 70%) compared to other models after tuning with our dataset. Notably, the performance improves significantly with the use of small context, achieving results comparable to those obtained with full context. This offers the potential for more efficient processing and reducing computational costs, allowing additional space to integrate various other types of context. In summary, our proposed instruction-tuning method can improve the model’s ability to both reuse dependencies and ensure the functional correctness of the output. Although self-refinement through multi-round debugging (Section 5.3) demonstrates greater effectiveness, instruction-tuning models only utilize single-turn generation, making them more efficient in practice. To facilitate open research in this domain, we will publicly release this dataset.
Model Full context Small context Pass@1 DIR Pass@1 DIR StarCoder 28.08 58.67 25.49 56.67 StarCoderDepIT 29.43 69.80 28.73 71.48 StarCoder2 27.77 60.57 23.27 53.49 StarCoder2DepIT 28.45 69.76 27.27 73.98 CodeLlama 38.65 62.26 35.66 62.67 CodeLlamaDepIT 38.85 68.89 36.93 73.19
@@ 6 Conclusion
We introduce RepoExec, a novel Python code generation benchmark at the repository level with executable capabilities, designed to evaluate the alignment of generated code with developer intent and its correctness. Our experiments reveal that while pretrained language models (LLMs) excel in functional correctness, instruction-tuned models demonstrate proficiency in utilizing dependencies and debugging. However, existing models struggle to effectively reuse provided dependencies, potentially leading to technical debt and code smell issues. We provide a comprehensive evaluation of how different LLMs leverage code dependencies when generating code, and our findings show that the size of the context significantly impacts the final results.
We also introduce an instruction-tuning dataset that enhances dependency invocation accuracy and output correctness, even with limited context. This approach allows for the integration of additional context types and mitigates large token length constraints. Our contributions establish a foundation for future research in code generation, providing valuable evaluation techniques and insights to drive the development of more capable and reliable models.
@@ 7 Limitations
RepoExec is constructed for evaluating models based on existing works. However, this approach can potentially lead to data leakage, especially when modern models are trained on similar datasets. If the benchmark relies heavily on known works, there’s a risk that the model may inadvertently learn from specific patterns or features present in those works, compromising its generalization ability. Despite this concern, experimental results demonstrate that even modern models struggle to handle the challenges posed by RepoExec, indicating that the benchmark remains a valuable tool for assessing model performance. However, employing new repositories can help mitigate these phenomena.
In this work, we only consider one level of dependency context, specifically the dependencies directly called from the target function. While this simplification facilitates manageable analysis and model development, it may not fully capture the valuable context necessary for leveraging models effectively. However, incorporating deep dependencies could significantly extend the input length, posing challenges in managing long context inputs and potentially exceeding the maximum input length. Our approach has demonstrated promising outcomes with the short context version, creating opportunities for integrating additional input context. Future research could explore incorporating multiple levels of dependencies, creating a more comprehensive graph that includes transitive dependencies, indirect calls, and broader contextual information. By doing so, we could enhance the model’s understanding of code interactions and improve its ability to handle intricate software execution scenarios.
@@
Several code generation benchmarks utilize match-based metrics like Edit similarity (ES), BLEU, and CodeBLEU for evaluation [Ding et al., 2024, Shrivastava et al., 2023, Ding et al., 2023, Liao et al., 2023, Yin et al., 2018, Iyer et al., 2018]. These metrics are straightforward to apply and may exhibit a strong correlation with execution metrics such as Pass@k. However, they cannot accurately measure functional correctness. For instance, comparing two Python code snippets where the only difference is the ":" character could result in a high ES and BLEU score. Nevertheless, one snippet may contain a syntax error, highlighting a limitation in these metrics for assessing true functionality.
@@ Appendix B Coverage Enhancement Effectiveness
Weak unit tests may inadvertently allow incorrect implementations to be determined as correct. Even with human-written tests, the overlooked coverage rates lead to evaluations that are incomplete and potentially misleading Liu et al. [2023a]. We present evidence supporting this argument, underscoring the limitations of prior work on code generation within repository-level contexts. As depicted in Figure 5, enhancing the number of test cases and coverage rates leads to a significant increase in the identification of incorrect generated solutions, causing the Pass@1 score to drop markedly (by over 5%). We investigated several solutions and found that most of the generated results did not fully utilize the given context (considered as human-provided). Instead, they primarily focused on addressing the problem described in the given natural language description. This indirectly overlooks the developer’s intentions, such as testing edge or corner cases, highlighting the limitations in following and understanding the provided intent and dependency context in these models. In summary, these findings underscore the effectiveness and importance of maintaining high-quality test cases for evaluation purposes.
@@ D.1 BasePrompt
Example of Short Context import base64 import random import unicodedata import zlib from typing import Union from uuid import uuid4 from ._regex import * from .errors import InvalidInputError from .validation import is_snake_case, is_full_string, is_camel_case, is_integer, is_string CAMEL_CASE_REPLACE_RE = re.compile(r’([a-z]|[A-Z]+)(?=[A-Z])’) class InvalidInputError(TypeError): def __init__(self, input_data: Any): def is_string(obj: Any) -> bool: def is_camel_case(input_string: Any) -> bool: def camel_case_to_snake(input_string, separator=’_’): """ Convert a camel case string into a snake case one. (The original string is returned if is not a valid camel case string) *Example:* $>$ camel_case_to_snake(’ThisIsACamelStringTest’) # returns ’this_is_a_camel_case_string_test’ :param input_string: String to convert. :type input_string: str :param separator: Sign to use as separator. :type separator: str :return: Converted string. """
@@ D.2 InstructPrompt
Example of Prompt 1 for Short Context ### Instruction: Write a Python function camel_case_to_snake(input_string, separator=’_’)‘ to solve the following problem: """ Convert a camel case string into a snake case one. (The original string is returned if is not a valid camel case string) *Example:* $>$ camel_case_to_snake(’ThisIsACamelStringTest’) # returns this_is_a_camel_case_string_test :param input_string: String to convert. :type input_string: str :param separator: Sign to use as separator. :type separator: str :return: Converted string. """ ### Response: import base64 import random import unicodedata import zlib from typing import Union from uuid import uuid4 from ._regex import * from .errors import InvalidInputError from .validation import is_snake_case, is_full_string, is_camel_case, is_integer, is_string CAMEL_CASE_REPLACE_RE = re.compile(r’([a-z]|[A-Z]+)(?=[A-Z])’) class InvalidInputError(TypeError): def __init__(self, input_data: Any): def is_string(obj: Any) -> bool: def is_camel_case(input_string: Any) -> bool: def camel_case_to_snake(input_string, separator=’_’): """ Convert a camel case string into a snake case one. (The original string is returned if is not a valid camel case string) *Example:* $>$ camel_case_to_snake(’ThisIsACamelStringTest’) # returns this_is_a_camel_case_string_test :param input_string: String to convert. :type input_string: str :param separator: Sign to use as separator. :type separator: str :return: Converted string. """
@@
Example of Prompt 2 for Short Context ### Instruction import base64 import random import unicodedata import zlib from typing import Union from uuid import uuid4 from ._regex import * from .errors import InvalidInputError from .validation import is_snake_case, is_full_string, is_camel_case, is_integer, is_string CAMEL_CASE_REPLACE_RE = re.compile(r’([a-z]|[A-Z]+)(?=[A-Z])’) class InvalidInputError(TypeError): def __init__(self, input_data: Any): def is_string(obj: Any) -> bool: def is_camel_case(input_string: Any) -> bool: The provided code snippet includes necessary dependencies for implementing the camel_case_to_snake function. Write a Python function camel_case_to_snake(input_string, separator=’_’)‘ to solve the following problem: """ Convert a camel case string into a snake case one. (The original string is returned if is not a valid camel case string) *Example:* $>$ camel_case_to_snake(’ThisIsACamelStringTest’) # returns this_is_a_camel_case_string_test :param input_string: String to convert. :type input_string: str :param separator: Sign to use as separator. :type separator: str :return: Converted string. """ ### Response: def camel_case_to_snake(input_string, separator=’_’): """ Convert a camel case string into a snake case one. (The original string is returned if is not a valid camel case string) *Example:* $>$ camel_case_to_snake(’ThisIsACamelStringTest’) # returns this_is_a_camel_case_string_test :param input_string: String to convert. :type input_string: str :param separator: Sign to use as separator. :type separator: str :return: Converted string. """
@@ Appendix E Studied LLMs: Supplemental results
Indeed, Figure 6 illustrates the proportion of generated functions that are empty for each LLM using BasePrompt. The findings indicate that utilizing Medium context results in a substantial number of empty functions. This may be due to the input format of the context when using BasePrompt, which can mislead the model into interpreting dependency functions as few-shot examples. In the Medium context, the function bodies of dependencies are removed, making their format identical to the target function prompt. This similarity can mislead the LMs, resulting in empty solutions. Particularly, Starcoder-2 is heavily impacted by this issue, as over 31% of its generated results are empty functions, revealing a significant weakness of the model. Meanwhile, short context effectively decreases the occurrence of empty function generation by the model and, in certain instances, improves models’ ability in dependency calls (e.g. CodeLlama-13b-Python, WizardCoder-Python-13B-V1.0, and Mixtral-8x7B-Instruct-v0.1 in Table 3). We believe that the following reasons could contribute to this observation. Firstly, employing short context reduces the input token count, preventing truncation when exceeding the maximum length limit, thus allowing uninterrupted solution generation by the model. This reduced context enables models to concentrate exclusively on dependency signatures, thereby enhancing the probability that generated solutions effectively utilize these dependency token names. Moreover, function names hold substantial semantic value by delineating the function’s purpose. Many studies have underscored that code summarization heavily relies on extracting information from function names [Haldar and Hockenmaier, 2024, Mondal et al., 2023, Sontakke et al., 2022]. Therefore, this concise representation of dependencies has the potential to improve how models utilize dependencies in generating code.
Additionally, Table LABEL:tab:examples_gen presents examples that support our findings in Section 5.2. For instance, in the first example, we observe that the instruction-tuned model can effectively utilize the given dependencies to manage edge cases, whereas the pretrained model fails to do so. This supports our initial findings. Meanwhile, the second and third examples demonstrate that pretrained models often reimplement or devise workarounds instead of leveraging the available context. Besides, in the third example, the instruction-tuned model correctly identifies the relevant case but fails to generate the correct solution. This may suggest that hallucinations complicate the outputs generated by instruction-tuned models.

### v3

@@ 6.2 Generation with Multi-round Debugging
In this section, we examine the models’ self-refinement capabilities in enhancing generation performance. We provide the models with error output logs and ask them to fix the errors. We experiment with WizardCoder, GPT3.5, and CodeLlama-13b-Python. The number of rounds to debug is set to 3 and the input template is presented in Appendix F. In this experiment, we employ a greedy search algorithm to generate only a single output.
@@ 6.3 Instruction-tuning with Code Dependencies
While the multi-round debugging experiment has demonstrated effectiveness in leveraging given dependencies to provide correct solutions (Section 6.2), it requires a strong model to generate good test cases and can be time-consuming due to the repeated generation and execution of code and test cases. To address these challenges, we propose an instruction-tuning dataset for fine-tuning base LLMs. We collected the 1,555 most-starred repositories from 2018 onward and extracted functions with their corresponding dependencies, following the procedure outlined in Section 4.2. We obtained 154,818 functions, of which 57,746 samples include docstrings. From this set, we utilized 50K samples with docstrings and applied instruction prompts, while 80K samples adhered to the raw code format (Full context). Recognizing the potential of the Small context format, we allocated the remaining 20K samples to follow this structure. We fine-tuned Phi-2, StarCoder, StarCoder2, and CodeLlama-13b-Python models for 5 epochs with LoRa Hu et al. (2021) and used 10% of the training data as the validation set to select the best checkpoint.
Table 5 illustrates the efficacy of our training data. All 4 models demonstrate improvements in both Pass@1 and DIR after instruction tuning. Specifically, there is a slight increase of around 1% in Pass@k for all models, while DIR shows a significant improvement, reaching the highest scores (over 70%) compared to other models after tuning with our dataset. Notably, the performance improves significantly with the use of small context, achieving results comparable to those obtained with full context. This offers the potential for more efficient processing and reducing computational costs, allowing additional space to integrate various other types of context. In summary, our proposed instruction-tuning method can improve the model’s ability to both reuse dependencies and ensure the functional correctness of the output. Although self-refinement through multi-round debugging (Section 6.2) demonstrates greater effectiveness, instruction-tuning models only utilize single-turn generation, making them more efficient in practice. To facilitate open research in this domain, we will publicly release this dataset.
Model Full context Small context Pass@1 DIR Pass@1 DIR phi-2 19.04 48.22 14.82 44.54 phi-2DepIT 20.20 61.66 20.31 70.30 StarCoder 28.08 58.67 25.49 56.67 StarCoderDepIT 29.43 69.80 28.73 71.48 StarCoder2 27.77 60.57 23.27 53.49 StarCoder2DepIT 28.45 69.76 27.27 73.98 CodeLlama 38.65 62.26 35.66 62.67 CodeLlamaDepIT 38.85 68.89 36.93 73.19
@@ 7 Conclusion
We introduce an evaluation approach for repository-level code generation that rethinks the limitations of previous methods by assessing not only functional correctness but also dependency utilization to ensure the quality of the generated code. By presenting RepoExec, a novel Python code generation benchmark at the repository level with executable capabilities, designed to evaluate the alignment of generated code with developer intent and its correctness. Our experiments reveal that while pretrained language models (LLMs) excel in functional correctness, instruction-tuned models demonstrate proficiency in utilizing dependencies and debugging. However, existing models struggle to effectively reuse provided dependencies, potentially leading to technical debt and code smell issues. We provide a comprehensive evaluation of how different LLMs leverage code dependencies when generating code, and our findings show that the size of the context significantly impacts the final results.
We also introduce an instruction-tuning dataset that enhances dependency invocation accuracy and output correctness, even with limited context. This approach allows for the integration of additional context types and mitigates large token length constraints. Our contributions establish a foundation for future research in code generation, providing valuable evaluation techniques to drive the development of more capable and reliable models.
@@ 8 Limitations
RepoExec is constructed for evaluating models based on existing works. However, this approach can potentially lead to data leakage, especially when modern models are trained on similar datasets. If the benchmark relies heavily on known works, there’s a risk that the model may inadvertently learn from specific patterns or features present in those works, compromising its generalization ability. Despite this concern, experimental results demonstrate that even modern models struggle to handle the challenges posed by RepoExec, indicating that the benchmark remains a valuable tool for assessing model performance. Besides, several models, including StarCoder2 and DeepSeek-Coder, have been pretrained using repository-level context. However, these models typically concatenate the contents of multiple files in a repository without filtering out irrelevant information or considering the selection of dependencies. This helps our dataset distinguish from the pretraining datasets of these models, thereby helping to mitigate the data leakage issue. However, employing new repositories can help mitigate these phenomena.
In this work, we only consider one level of dependency context, specifically the dependencies directly called from the target function. While this simplification facilitates manageable analysis and model development, it may not fully capture the valuable context necessary for leveraging models effectively. However, incorporating deep dependencies could significantly extend the input length, posing challenges in managing long context inputs and potentially exceeding the maximum input length. Our approach has demonstrated promising outcomes with the small-size context version, creating opportunities for integrating additional input context. Future research could explore incorporating multiple levels of dependencies, creating a more comprehensive graph that includes transitive dependencies, indirect calls, and broader contextual information. By doing so, we could enhance the model’s understanding of code interactions and improve its ability to handle intricate software execution scenarios.
@@
Several code generation benchmarks utilize match-based metrics like Edit similarity (ES), BLEU, and CodeBLEU for evaluation (Ding et al., 2024; Shrivastava et al., 2023; Ding et al., 2023; Liao et al., 2023; Yin et al., 2018; Iyer et al., 2018). These metrics are straightforward to apply and may exhibit a strong correlation with execution metrics such as Pass@k. However, they cannot accurately measure functional correctness. For instance, comparing two Python code snippets where the only difference is the ":" character could result in a good ES and BLEU score. Nevertheless, one snippet may contain a syntax error, highlighting a limitation in these metrics for assessing true functionality.
@@ Appendix B Coverage Enhancement Effectiveness
Weak unit tests may inadvertently allow incorrect implementations to be determined as correct. Even with human-written tests, the overlooked coverage rates lead to evaluations that are incomplete and potentially misleading Liu et al. (2023a). We present evidence supporting this argument, underscoring the limitations of prior work on code generation within repository-level contexts. As depicted in Figure 5, enhancing the number of test cases and coverage rates leads to a significant increase in the identification of incorrect generated solutions, causing the Pass@1 score to drop markedly (by over 5%). We investigated several solutions and found that most of the generated results did not fully utilize the given context (considered as human-provided). Instead, they primarily focused on addressing the problem described in the given natural language description. This indirectly overlooks the developer’s intentions, such as testing edge or corner cases, highlighting the limitations in following and understanding the provided intent and dependency context in these models. In summary, these findings underscore the effectiveness and importance of maintaining high-quality test cases for evaluation purposes.
@@ D.1 BasePrompt
Example of Small Context import base64 import random import unicodedata import zlib from typing import Union from uuid import uuid4 from ._regex import * from .errors import InvalidInputError from .validation import is_snake_case, is_full_string, is_camel_case, is_integer, is_string CAMEL_CASE_REPLACE_RE = re.compile(r’([a-z]|[A-Z]+)(?=[A-Z])’) class InvalidInputError(TypeError): def __init__(self, input_data: Any): def is_string(obj: Any) -> bool: def is_camel_case(input_string: Any) -> bool: def camel_case_to_snake(input_string, separator=’_’): """ Convert a camel case string into a snake case one. (The original string is returned if is not a valid camel case string) *Example:* $>$ camel_case_to_snake(’ThisIsACamelStringTest’) # returns ’this_is_a_camel_case_string_test’ :param input_string: String to convert. :type input_string: str :param separator: Sign to use as separator. :type separator: str :return: Converted string. """
@@ D.2 InstructPrompt
Example of Prompt 1 for Small Context ### Instruction: Write a Python function camel_case_to_snake(input_string, separator=’_’)‘ to solve the following problem: """ Convert a camel case string into a snake case one. (The original string is returned if is not a valid camel case string) *Example:* $>$ camel_case_to_snake(’ThisIsACamelStringTest’) # returns this_is_a_camel_case_string_test :param input_string: String to convert. :type input_string: str :param separator: Sign to use as separator. :type separator: str :return: Converted string. """ ### Response: import base64 import random import unicodedata import zlib from typing import Union from uuid import uuid4 from ._regex import * from .errors import InvalidInputError from .validation import is_snake_case, is_full_string, is_camel_case, is_integer, is_string CAMEL_CASE_REPLACE_RE = re.compile(r’([a-z]|[A-Z]+)(?=[A-Z])’) class InvalidInputError(TypeError): def __init__(self, input_data: Any): def is_string(obj: Any) -> bool: def is_camel_case(input_string: Any) -> bool: def camel_case_to_snake(input_string, separator=’_’): """ Convert a camel case string into a snake case one. (The original string is returned if is not a valid camel case string) *Example:* $>$ camel_case_to_snake(’ThisIsACamelStringTest’) # returns this_is_a_camel_case_string_test :param input_string: String to convert. :type input_string: str :param separator: Sign to use as separator. :type separator: str :return: Converted string. """
@@
Example of Prompt 2 for Small Context ### Instruction import base64 import random import unicodedata import zlib from typing import Union from uuid import uuid4 from ._regex import * from .errors import InvalidInputError from .validation import is_snake_case, is_full_string, is_camel_case, is_integer, is_string CAMEL_CASE_REPLACE_RE = re.compile(r’([a-z]|[A-Z]+)(?=[A-Z])’) class InvalidInputError(TypeError): def __init__(self, input_data: Any): def is_string(obj: Any) -> bool: def is_camel_case(input_string: Any) -> bool: The provided code snippet includes necessary dependencies for implementing the camel_case_to_snake function. Write a Python function camel_case_to_snake(input_string, separator=’_’)‘ to solve the following problem: """ Convert a camel case string into a snake case one. (The original string is returned if is not a valid camel case string) *Example:* $>$ camel_case_to_snake(’ThisIsACamelStringTest’) # returns this_is_a_camel_case_string_test :param input_string: String to convert. :type input_string: str :param separator: Sign to use as separator. :type separator: str :return: Converted string. """ ### Response: def camel_case_to_snake(input_string, separator=’_’): """ Convert a camel case string into a snake case one. (The original string is returned if is not a valid camel case string) *Example:* $>$ camel_case_to_snake(’ThisIsACamelStringTest’) # returns this_is_a_camel_case_string_test :param input_string: String to convert. :type input_string: str :param separator: Sign to use as separator. :type separator: str :return: Converted string. """
@@ Appendix E Studied LLMs: Supplemental results
Indeed, Figure 6 illustrates the proportion of generated functions that are empty for each LLM using BasePrompt. The findings indicate that utilizing Medium context results in a substantial number of empty functions. This may be due to the input format of the context when using BasePrompt, which can mislead the model into interpreting dependency functions as few-shot examples. In the Medium context, the function bodies of dependencies are removed, making their format identical to the target function prompt. This similarity can mislead the LMs, resulting in empty solutions. Particularly, Starcoder-2 is heavily impacted by this issue, as over 31% of its generated results are empty functions, revealing a significant weakness of the model. Meanwhile, small context effectively decreases the occurrence of empty function generation by the model and, in certain instances, improves models’ ability in dependency calls (e.g. CodeLlama-13b-Python, WizardCoder-Python-13B-V1.0, and Mixtral-8x7B-Instruct-v0.1 in Table 3). We believe that the following reasons could contribute to this observation. Firstly, employing small context reduces the input token count, preventing truncation when exceeding the maximum length limit, thus allowing uninterrupted solution generation by the model. This reduced context enables models to concentrate exclusively on dependency signatures, thereby enhancing the probability that generated solutions effectively utilize these dependency token names. Moreover, function names hold substantial semantic value by delineating the function’s purpose. Many studies have underscored that code summarization heavily relies on extracting information from function names (Haldar and Hockenmaier, 2024; Mondal et al., 2023; Sontakke et al., 2022). Therefore, this concise representation of dependencies has the potential to improve how models utilize dependencies in generating code.
Additionally, Table LABEL:tab:examples_gen presents examples that support our findings in Section 6.1. For instance, in the first example, we observe that the instruction-tuned model can effectively utilize the given dependencies to manage edge cases, whereas the pretrained model fails to do so. This supports our initial findings. Meanwhile, the second and third examples demonstrate that pretrained models often reimplement or devise workarounds instead of leveraging the available context. Besides, in the second example, the instruction-tuned model correctly identifies the relevant case but fails to generate the correct solution. This may suggest that hallucinations complicate the outputs generated by instruction-tuned models.
@@ Appendix F Multi-round Debugging
Appendix G Context length analysis
Long-context models can enhance the ability to comprehend and select relevant context from lengthy inputs to effectively solve the required task. In this section, we examine how the context length supported by each model affects their performance in RepoExec. Figure 9 demonstrates the relationship between support context length, model size, and model family in relation to pass@k and DIR scores. The size of the dots indicates the model size, while the color represents the model family.
Figure 9: Correlation of context length to the model performance on RepoExec.
We observe that context length has a weak correlation with model performance on our dataset. In contrast, model size (scaling law) and model family, which encompass different training methods (pretraining or instruction tuning), training datasets, and architectures, show a more significant impact. The weak correlation with context length can be explained by our approach’s ability to already capture relevant information (e.g. dependency) for each data sample, resulting in the pruning of context length for practical usage (363 tokens on average shown in Table 2). Meanwhile, models that support long contexts are often trained on data containing a mix of relevant and irrelevant information and are evaluated on their ability to retrieve the correct context in a needle-in-a-haystack scenario (Roziere et al., 2023; Ivgi et al., 2023; Liu et al., 2024). Therefore, models with varying context lengths might show a weak correlation to performance in our scenario.
Appendix H Dependency extraction tool usage
We present pydepcall, a Python library designed to extract function dependencies from any repository. We provide a brief overview of its usage in the following code snippet.
pydepcall usage from pydepcall import Extractor # If you want to extract all module files in the repository reposrc = YOUR_LOCAL_PATH_OF_REPO extractor = Extractor(reposrc) output = extractor.extract() # If you want to extract a specific module file in the repository reposrc = YOUR_LOCAL_PATH_OF_REPO module_file = YOUR_LOCAL_PATH_OF_FILE_IN_REPO extractor = Extractor(reposrc, module_file) output = extractor.extract()